# A multidimensional measurement of photorealistic avatar quality of experience

### v2

@@ 4.3. Reproducibility
To test the reproducibility from run to run, for each template we repeated our crowdsourcing test five times with a mutually exclusive group of workers, on separate days on Amazon Mechanical Turk. On average we have collected 22.8 votes per video clip using Template A and 22.4 votes using Template B in each run. We calculated the Mean Opinion Scores (MOS) per clip and per condition (avatar variation) and show the correlation between different runs and scales in Tables 7 and 8. The results show a strong correlation between different runs at the clip (average PCC=0.968) and condition levels (average PCC=0.987) and confirm the survey provides reproducible scores.
@@ 5.2. Correlation analysis for Template A
We observed a strong correlation between the dimensions from Template A (Table 9), on average PCC=0.996. This was surprisingly high, which motivated doing the three versions of the survey in Section 4.1 which all exhibited the same high correlation. However, when we consider only low realism avatars (realism $\leq$ 2) then the dimensions are not all highly correlated as shown in Table 11. But for avatars with realism ¿ 2, the dimensions are very highly correlated with an average PCC=0.999 as shown in Table 10. One important implication of this high correlation is that avatars that are not as realistic as real video will have lower trust, comfortableness using, comfortableness interacting with, appropriateness for work, formality, and affinity, and higher creepiness compared to real video.
@@
Table 10. PCC across dimensions when realism ¿ 2
@@ 5.4. Correlation with objective metrics
We evaluated the effectiveness of objective metrics in representing subjective scores. In this domain, PSNR (Gonzalez and Woods, 2006), SSIM (Wang et al., 2004), and LPIPS (Zhang et al., 2018) are typically reported in publications. We also analyze FID (Heusel et al., 2017) and FVD (Unterthiner et al., 2019), which are less commonly reported. We utilized eight models from the study reported in Section 4.3, where the original video was available (overall 31 clips). We removed the background of the person from the original videos and transformed the videos to match each avatar representation using face landmarks. Finally, we calculated the objective metrics for each video clip and aggregated them for the avatar models. Table 12 reports the Pearson’s correlation coefficient and Kendall’s Tau-b rank correlation coefficient [^3] We report Kendall’s Tau-b as it is more robust to outliers and non-normal distributions compare to Spearman’s rank correlation (Croux and Dehon, 2010) making it more suitable for small sample sizes. (Kendall, 1938) between the objective metrics and the subjective scores from Template A and Template B. The results show a weak correlation between the objective metrics and the subjective scores from Template A, and a moderate correlation with scores from Template B for PSNR, SSIM, and LPIPS. This highlights the necessity of using subjective scores for accessing the quality of experience of photorealistic avatars.
@@ 6. Conclusions and future work
We have provided a test framework to measure the quality of experience of photorealistic avatars, and shown it to be accurate and reproducible (RQ3). We have shown the commonly used objective metrics (PSNR, SSIM, LPIPS, FIV, FVD) are weakly correlated to nine of the subjective metrics we define, and moderately correlated to emotion accuracy (RQ1). For avatars with a realism $>$ 2 MOS we have shown in Table 10 that all of the no-reference metrics (Template A) are highly correlated to each other, which means that for Template A only one metric needs to be measured (e.g., realism), while for Template B both metrics need to be measured; this leads to dimensionality reduction of 10 to 3 for the survey (RQ2). We have shown in the context of photorealistic avatars in telecommunication scenarios that the uncanny valley effect does not exist, which implies photorealistic avatars should be made as realistic as possible to maximize the quality of experience of the avatar (RQ4). Finally, we show that avatars that are not as realistic as real video will have lower trust, comfortableness using, comfortableness interacting with, appropriateness for work, formality, and affinity, and higher creepiness compared to real video (RQ5).
@@
  • During the development of collaboration systems with photorealistic avatars regular subjective tests should be conducted as we have shown the objective metrics are not well correlated to the subjective metrics in this study.
One limitation of this study is the diversity of the avatars studied. We included 20 avatars which included 67 males, 21 females, 51 Caucasians, 27 Asians, and 10 Blacks. Ideally, we would have included more avatars with an equal number of males and females and a greater number of and more equality distributed races represented.
Finally, the use of photorealistic avatars in telecommunication systems introduces potential privacy, trust, and ethical considerations, which are outside the scope of this paper and should be addressed by the community.
@@ Acknowledgements.
We thank [anonymous] for providing some avatars for this study.

### v3

@@ 4.3. Reproducibility
To test the reproducibility from run to run, for each template we repeated our crowdsourcing test five times with a mutually exclusive group of workers, on separate days on Amazon Mechanical Turk. On average we have collected 22.8 votes per video clip using Template A and 22.4 votes using Template B in each run. We calculated the MOS per clip and per condition (avatar variation) and show the correlation between different runs and scales in Tables 7 and 8. The results show a strong correlation between different runs at the clip (average PCC=0.968) and condition levels (average PCC=0.987) and confirm the survey provides reproducible scores.
@@ 5.1. Results
Template A Template B
Comfortable Comfortable Emotion
Appropriate 1.000 0.997 1.000 0.998 1.000 0.999 0.997 1.000 0.650 0.390
Comfortable interacting 0.997 1.000 0.995 0.993 0.996 0.997 0.996 0.996 0.651 0.407
Comfortable using 1.000 0.995 1.000 0.998 0.999 0.998 0.997 1.000 0.653 0.398
Formal 0.998 0.993 0.998 1.000 0.997 0.999 0.997 0.998 0.659 0.406
Affinity 1.000 0.996 0.999 0.997 1.000 0.998 0.995 0.999 0.636 0.372
Not Creepy 0.999 0.997 0.998 0.999 0.998 1.000 0.997 0.998 0.646 0.385
Realistic 0.997 0.996 0.997 0.997 0.995 0.997 1.000 0.997 0.669 0.454
Trust 1.000 0.996 1.000 0.998 0.999 0.998 0.997 1.000 0.656 0.404
Emotion accuracy 0.650 0.651 0.653 0.659 0.636 0.646 0.669 0.656 1.000 0.596
Resemblance 0.390 0.407 0.398 0.406 0.372 0.385 0.454 0.404 0.596 1.000
@@ 5.2. Correlation analysis for Template A
We observed a strong correlation between the dimensions from Template A (Table 9), on average PCC=0.996. This was surprisingly high, which motivated doing the three versions of the survey in Section 4.1 which all exhibited the same high correlation. However, when we consider only low realism avatars (Realistic $\leq$ 2) then the dimensions are not all highly correlated as shown in Table 11. But for avatars with Realistic ¿ 2, the dimensions are very highly correlated with an average PCC=0.999 as shown in Table 10. One important implication of this high correlation is that avatars that are not as realistic as real video will have lower trust, comfortableness using, comfortableness interacting with, appropriateness for work, formality, and affinity, and higher creepiness compared to real video.
@@
Table 10. PCC across dimensions when realism ¿ 2.
Comfortable Comfortable
Appropriate interacting using Formal Affinity Not Creepy Realistic Trust
Appropriate 1.000 0.998 0.999 0.997 0.999 0.996 0.997 0.999
Comfortable interacting 0.998 1.000 0.998 0.996 0.998 0.998 0.997 0.997
Comfortable using 0.999 0.998 1.000 0.997 0.998 0.996 0.996 0.998
Formal 0.997 0.996 0.997 1.000 0.997 0.993 0.997 0.998
Affinity 0.999 0.998 0.998 0.997 1.000 0.996 0.996 0.998
Not Creepy 0.996 0.998 0.996 0.993 0.996 1.000 0.994 0.995
Realistic 0.997 0.997 0.996 0.997 0.996 0.994 1.000 0.997
Trust 0.999 0.997 0.998 0.998 0.998 0.995 0.997 1.000
@@
Comfortable Comfortable
Appropriate interacting using Formal Affinity Not Creepy Realistic Trust
Appropriate 1.000 0.824 0.909 0.704 0.885 0.655 0.763 0.894
Comfortable interacting 0.824 1.000 0.854 0.575 0.879 0.806 0.771 0.797
Comfortable using 0.909 0.854 1.000 0.693 0.908 0.681 0.763 0.879
Formal 0.704 0.575 0.693 1.000 0.707 0.384 0.670 0.706
Affinity 0.885 0.879 0.908 0.707 1.000 0.732 0.783 0.871
Not Creepy 0.655 0.806 0.681 0.384 0.732 1.000 0.549 0.600
Realistic 0.763 0.771 0.763 0.670 0.783 0.549 1.000 0.727
Trust 0.894 0.797 0.879 0.706 0.871 0.600 0.727 1.000
@@ 5.4. Correlation with objective metrics
We evaluated the effectiveness of objective metrics in representing subjective scores. In this domain, PSNR (Gonzalez and Woods, 2006), SSIM (Wang et al., 2004), and LPIPS (Zhang et al., 2018) are typically reported in publications. We also analyze FID (Heusel et al., 2017) and FVD (Unterthiner et al., 2019), which are less commonly reported. We utilized eight models from the study reported in Section 4.3, where the original video was available (overall 31 clips). We removed the background of the person from the original videos and transformed the videos to match each avatar representation using face landmarks (see Figure 9 for an example). Finally, we calculated the objective metrics for each video clip and aggregated them for the avatar models. Table 12 reports the Pearson’s correlation coefficient and Kendall’s Tau-b rank correlation coefficient [^3] We report Kendall’s Tau-b as it is more robust to outliers and non-normal distributions compare to Spearman’s rank correlation (Croux and Dehon, 2010) making it more suitable for small sample sizes. (Kendall, 1938) between the objective metrics and the subjective scores from Template A and Template B. The results show a weak correlation between the objective metrics and the subjective scores from Template A, and a moderate correlation with scores from Template B for PSNR, SSIM, and LPIPS. This highlights the necessity of using subjective scores for accessing the quality of experience of photorealistic avatars.
((a))
((b))
((c))
Figure 9. Modifying the original video to match the position of the avatar using face landmarks and removing the background before calculating objective metrics. (a) A frame from the original video, (b) the corresponding frame from the avatar, (c) an overlay of the modified original frame and the avatar’s frame.
@@ 6. Conclusions and future work
We have shown the commonly used objective metrics (PSNR, SSIM, LPIPS, FIV, FVD) are weakly correlated to nine of the subjective metrics we define, and moderately correlated to emotion accuracy (RQ1). Note that none of these metrics are trained to predict avatar realism. More sophisticated objective metrics may improve objective performance, especially if trained using a large subjective dataset of avatar realism (this is done in audio (Yi et al., 2022) and video (Majeedi et al., 2023) objective quality models).
For avatars with a realism $>$ 2 MOS we have shown in Table 10 that all of the no-reference metrics (Template A) are highly correlated to each other, which means that for Template A only one metric needs to be measured (e.g., realism), while for Template B both metrics need to be measured; this leads to dimensionality reduction of 10 to 3 for the survey (RQ2).
We have provided an open source crowdsourced test framework to measure the quality of experience of photorealistic avatars, and shown it to be accurate and reproducible (RQ3).
We have shown in the context of photorealistic avatars in telecommunication scenarios that the uncanny valley effect does not exist, which implies photorealistic avatars should be made as realistic as possible to maximize the quality of experience of the avatar (RQ4). Note that uncanny valley effects may still exist in other domains (e.g., gaming, metaverse) and with other visual distortions (e.g., motion artifacts or mismatches to expected expressions).
Finally, we show that avatars that are not as realistic as real video will have lower trust, comfortableness using, comfortableness interacting with, appropriateness for work, formality, and affinity, and higher creepiness compared to real video (RQ5). One implication is that to get the level trust, comfortableness using, comfortableness interacting with, appropriateness for work, formality, and affinity, and higher creepiness as real video, the avatar needs the same level of realism as real video, which is a very high bar not yet achieved by any of the avatars tested.
6.1. Extensions
@@
  • The test framework could be extended to 3D environments such as virtual reality (VR) headsets.
  • The uncanny valley effect can be further studied using attractiveness, eeriness, humanness, and spine-tingling.
6.2. Developer guidelines
@@
  • During the development of collaboration systems with photorealistic avatars regular subjective tests should be conducted as we have shown the objective metrics are not well correlated to the subjective metrics in this study. Not all ten dimensions need to be measured as we have shown only 3 are needed. This greatly simplifies the measurement process as well as reducing the number of dimensions that need to be optimized during development.
6.3. Limitations
There are several limitations for this study, including:
  • The number of avatars used in the study (N=19) would ideally be much larger (e.g., N=50). In addition, the avatars should ideally be all tested with identical material, which was only the case when the avatar model was available.
  • The diversity of the avatars studied: We used 19 avatar models to create 88 instances of avatars, which included 67 males, 21 females, 51 Caucasians, 27 Asians, and 10 Blacks. Ideally, we would have included more avatars with an equal number of males and females and a greater number of and more equally distributed races represented.
  • The test framework is a passive test and the results of actively using the avatars may be different.
  • The evaluations are done on 2D displays and the results may be different if tested on 3D displays or VR headsets.
  • The use of photorealistic avatars in telecommunication systems introduces potential privacy, trust, and ethical considerations, which are outside the scope of this paper and should be addressed by the community.
@@ Acknowledgements.
We thank various teams at Microsoft for providing some avatars for this study.